OpenStudy (anonymous):

A cheetah runs at a speed of 54 miles for every hour. If the distance traveled, in miles, is d and time, in hours, is t, which equation shows the relationship between d and t? d = 54t t = 54d d = 54 + t t = 54 + d @mathstudent55 @BloomLocke367 @triciaal @mathmath333 @kevindragonlord

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} \normalsize \text{use the formula}\hspace{.33em}\\~\\ speed=\dfrac{distance}{time}\hspace{.33em}\\~\\ \end{align}}\)

OpenStudy (mathmath333):

plug in \(\large \color{black}{\begin{align} speed=54\quad miles\\ distance=d\\ times =t \end{align}}\)

OpenStudy (mathstudent55):

Speed can be given in the units of miles per hour. Miles per hour means "miles divided by hour." Since miles are used for a distance, and hours are used for time, you can conclude that speed is equal to distance divided by time, or speed = distance/time, or speed = d/t If you multiply both sides of this equation by time, and you replace speed with 54 mph, whioch answer do you get?

OpenStudy (mathmath333):

after plugging in u get this \(\large \color{black}{\begin{align} 54=\dfrac{d}{t}\hspace{.33em}\\~\\ \end{align}}\) now multily \(t\) on both sides